# On computing and the complexity of computing higher-order $U$-statistics, exactly

### v1

@@ Abstract
Higher-order $U$-statistics abound in fields such as statistics, machine learning, and computer science, but are known to be highly time-consuming to compute in practice. Despite their widespread appearance, a comprehensive study of their computational complexity is surprisingly lacking. This paper aims to fill that gap by presenting several results related to the computational aspect of $U$-statistics. First, we derive a useful decomposition from an $m$-th order $U$-statistic to a linear combination of $V$-statistics with orders not exceeding $m$, which are generally more feasible to compute. Second, we explore the connection between exactly computing $V$-statistics and Einstein summation, a tool often used in computational mathematics, quantum computing, and quantum information sciences for accelerating tensor computations. Third, we provide an optimistic estimate of the time complexity for exactly computing $U$-statistics, based on the treewidth of a particular graph associated with the $U$-statistic kernel. The above ingredients lead to a new, much more runtime-efficient algorithm of exactly computing general higher-order $U$-statistics. We also wrap our new algorithm into an open-source Python package called u-stats. We demonstrate via three statistical applications that u-stats achieves impressive runtime performance compared to existing benchmarks. This paper aspires to achieve two goals: (1) to capture the interest of researchers in both statistics and other related areas further to advance the algorithmic development of $U$-statistics, and (2) to offer the package u-stats as a valuable tool for practitioners, making the implementation of methods based on higher-order $U$-statistics a more delightful experience.
Keywords: $U$-statistics, Statistical/Scientific Computing, Tensors, Einstein summation, Treewidth
@@ 1 Introduction
Higher-order $U$-statistics are prevalent in various disciplines, including statistics (robins2008higher; bonvini2022fast; bonvini2024doubly; neykov2021minimax; yao2018testing; breunig2024adaptive; shao2025u), machine learning (clemenccon2008ranking; clemenccon2013maximal; shen2025engression), random graphs (chatterjee2024higher), Bayesian networks and statistical physics (pearl1988probabilistic; koller2009probabilistic; shpitser2011efficient; li2020random), just to name a few. It is well-known in the statistical community that $U$-statistics are highly time-consuming to compute at large. In the worst-case, an $m$-th order $U$-statistic over a sample of size $n$ takes roughly $O(n^{m})$ steps to compute, because by definition (see Definition 1) one needs to sum over all $m$-out-of-$n$ permutations (resp. over $n^{m}$ terms). If $n=10^{4}$ and $m=7$, $n^{m}\approx 10^{28}$!
@@
More specifically, in causal inference, robins2008higher; robins2016technical; robins2017minimax developed a so-called “higher-order influence function (HOIF)” framework for constructing nearly rate-optimal estimators of average treatment effects under ignorability (hernan2023causal) under various structural assumptions, using higher-order $U$-statistics with order roughly of $\sqrt{\log n}$. Despite their appealing statistical properties, implementing such complex estimators remains challenging in practice and computational complexity is one important bottleneck. Here we quote a sentence from a recent graduate-level textbook on causal inference by wager2024causal (page 44) to highlight such a quandary:
@@
One challenge with the HOIF estimator of robins2017minimax, however, is that to date it has been challenging to implement in practical applications …
@@
This paper was initially motivated by (at least partially) addressing the computational aspect of this challenge, but we will present general-purpose theory and algorithm for computing general higher-order $U$-statistics beyond those for HOIF estimators. We will revisit HOIF estimators in Section 4.1.
@@
To tackle the difficulty of computing higher-order $U$-statistics, statisticians and computer scientists have long considered the so-called “incomplete $U$-statistics”, where one sums over only a random subset of all $m$-out-of-$n$ permutations (blom1976some; janson1984asymptotic; nowicki1988subgraph; shao2025u). Another natural approach is to choose a subsample of size $n^{\prime}<n$ from the whole sample and then compute a complete $U$-statistic from that subsample. Both strategies are akin to randomized or sketching algorithms in the theoretical computer science literature (mahoney2011randomized). The randomization approach certainly injects extra randomness in the final statistical results, which may result in inconsistent statistical conclusions across different randomly selected permutations/subsamples (guo2025rank). For a restricted class of higher-order $U$-statistics with decomposable and symmetric (or more precisely permutation-invariant) kernels, kong2018estimating cleverly recognize that exact computation can be accomplished by repeated matrix multiplications. A somewhat similar strategy has also been considered in zhang2025adaptive, in which an incomplete $U$-statistic with a special kernel structure is computed more efficiently via dynamic programming. For asymmetric kernels, their strategy essentially computes an incomplete $U$-statistic instead. All the above approaches may suffer from efficiency loss by not using all the available information.
The literature on (the complexity of) exactly computing higher-order $U$-statistics is rather scarce. This is hardly surprising because the exact time complexity of computing higher-order $U$-statistics seems to simply equal the total number of summations. However, as we will demonstrate in the paper, exactly computing $U$-statistics can be improved compared to a brute-force summation over all summands (using for-loops) when the kernel enjoys a particular decomposable structure specified in Definition 9 later. Interestingly, as we will exhibit in Section 4, such a structure holds in many statistical applications. It will also be clear later that it is helpful to bring in tools such as the Einstein summation from the computational mathematics community to obtain improved algorithms of computing $U$-statistics by leveraging such a structure. It is worth noting that Section 3 of liu2024hypothesis considered to use di-graphs to represent the structure of a class of $U$-statistics, a strategy conceptually similar to ours. Yet our new algorithm is designed to cover general higher-order $U$-statistics.
@@
The main contributions of our paper can be summarized as below.
  • We design a new algorithmic framework for exactly computing higher-order $U$-statistics and another closely related objects called $V$-statistics by leveraging the Einstein summation operation and related concepts from combinatorics and graph theory. Thanks to the connection with graph theory, we can provide an optimistic estimate of the time complexity of exactly computing $U$-statistics (see Section 3.3 for details). Although the idea is conceptually simple, to the best of our knowledge, it has not been thoroughly explored in the existing statistical literature.
  • Despite the ubiquity of $U$-statistics (see later this section), a dedicated, general-purpose software package for computing them seems to be lacking. To fill in this void, we developed a Python package called u-stats, which implemented the aforementioned algorithm. We demonstrate that the runtime of u-stats achieves the state-of-the-art performance in various examples, including computing HOIF estimators, motif counts of networks, and the distance covariance in the independence testing problem. In fact, as will be clear in Section 4, u-stats can dramatically improve the runtime compared to popular benchmarks. u-stats currently supports both CPU and GPU infrastructure. Having access to GPU can further improve the runtime because our algorithm heavily utilizes tensor operations such as the Einstein summation. We envision that the package u-stats can be a valuable tool for practitioners.
Remark 1.
We focus on scalar-valued $U$-statistics in this paper, but we plan to upgrade our package u-stats to handle vector- or matrix-valued cases in its next version. In this paper, one floating point operation is considered as the basic computation unit.
@@
In this part, we provide an incomplete survey on the related literature of $U$-statistics, and the relevant techniques/tools for developing our new algorithm.
@@
$U$-statistics are frequently found in statistics and machine learning. For example, in certain machine learning tasks (clemenccon2008ranking; clemenccon2013maximal; shen2025engression), the empirical training risk can be written as a $U$-statistic instead of a sample mean. Various (conditional) dependence measures can also be represented as $U$-statistics (yao2018testing; neykov2021minimax; chatterjee2021new; azadkia2021simple; shao2025u; zhang2025adaptive); see Section 4.3 for an example. As mentioned before, in causal inference, HOIF has been shown to be a theoretically appealing framework for constructing nearly rate-optimal estimators for various causal effect parameters. Estimators based on HOIF are higher-order $U$-statistics; see Section 4.1 for concrete forms. As mentioned, due to the computational difficulty, HOIF estimators have not been comprehensively applied in real-life applications.
@@
Higher-order $U$-statistics also appear very often in statistical inference over networks, often modeled as random graphs (chatterjee2024higher; jin2025counting). For instance, chatterjee2024higher estimate motif counts (such as triangles, 4-cliques, and etc.) by higher-order $U$-statistics and use these estimated motif counts to test certain properties of the underlying network. Motif counts also have widespread applications in bioinformatics (wernicke2006fanmod). As will be shown in Section 4.2, using the state-of-the-art software package igraph, for a moderately dense network, computing motif counts can be quite time consuming, necessitating the development of new algorithmic tools for $U$-statistics.
@@
Since computing UU/$V$-statistic is closely related to numerical integration using Monte Carlo sampling for which tensor operations are quite useful, our work is inspired by related techniques developed in the computational mathematics and physics communities. In particular, our algorithm utilizes the Einstein summation operation, an important tool for tensor computations. For example, in quantum computing/information, tensor networks are often used to describe many-body quantum systems and a very useful operation called tensor contraction is implemented by the Einstein summation in practice (markov2008simulating; stoudenmire2018learning). As will be seen later, our proposed algorithm is similar to tensor contraction in spirit.
Probabilistic graphical models and statistical physics
Our work is also related to the literature on probabilistic graphical models (pearl1988probabilistic; shpitser2011efficient), in which a common problem is to compute certain marginal distributions from the joint distribution, (nested) factorized according to Bayesian network. It can be reduced to computing $U$-statistics with kernels satisfying the decomposable structure to be defined later. Similarly, in both classical (edwards1975theory; li2020random) and quantum many-body systems (bernaschi2024quantum), the corresponding log-Gibbs measure can often be written as a multi-linear sum over all particles. The summation is taken over all possible configurations of each particle. Thus the log-Gibbs measure can be viewed as (a linear combination of) higher-order $U$-statistics.
@@
In this section, we first collect some frequently used notation. We adopt the following convention: we use the demarcation $\{\}$ for a set, representing an unordered collection of distinct elements; whereas we use the demarcation $()$ for a tuple, meaning an ordered collection of elements which may contain duplicates. Given a positive integer $m$, we let $m]\coloneqq\{1,2,\ldots,m\}$. To reduce clutter, we write $\bar{i}_{m}\equiv(i_{1},\cdots,i_{m})$ for short. $\mathbb{N}$ denotes the set of natural numbers including zero, while $\mathbb{N}^{+}$ denotes the set of all positive integers. Given a tuple $\bm{s}$, we let $\mathsf{set}(\bm{s})\coloneqq\{s_{k}\mid k\in[m]\}$ denote the set containing all elements appear in $\bm{s}$. Given a non-empty and finite set $S$ and a positive integer $m$, all $m$-permutations of elements in $S$ is denoted by $\mathsf{Perm}(S,m)$. All partitions of $S$ is denoted by $\Pi_{S}$. If $S$ happens to be integer set $[m]$, $\Pi_{S}$ is simplified to $\Pi_{m}$. Given sets $S_{1},\cdots,S_{m}$, $\times_{i=1}^{m}S_{i}$ denotes their Cartesian product. When $S_{1}=\cdots=S_{m}=S$, we simply write $S^{m}\equiv\times_{i=1}^{m}S_{i}$. $S^{*}\coloneqq\cup_{m\in\mathbb{N}^{+}}S^{m}$ denotes all tuples possibly generated by $S$. Given a tuple $\bm{s}\in S^{m}$ and a tuple $\alpha\in[m]^{*}$, we let $\bm{s}[\alpha]\coloneqq(s_{\alpha_{k}})_{k=1}^{|\alpha|}$.
@@
Terminologies from graph theory will be frequently encountered. Throughout this paper, all graphs are assumed to be finite and undirected without self-loops, but are allowed to have isolated vertices, unless stated otherwise. A graph $G$ is denoted by a pair $(V,E)$, where $V\equiv V(G)$ is the finite vertex set, and $E\equiv E(G)$ is the edge set of $G$. Two vertices $u$ and $v$ are said to be adjacent if $\{u,v\}\in E$. For any vertex $v\in V$, its neighbor set is denoted as $N_{G}(v)\coloneqq\{u\in V\mid\{u,v\}\in E\}$, and the degree of $v$ is defined as $\deg(v,G)\coloneqq|N_{G}(v)|$. A graph $G^{\prime}$ is a subgraph of graph $G$, if $V(G^{\prime})\subseteq V(G)$ and $E(G^{\prime})\subseteq E(G)$. For any subset $V^{\prime}\subseteq V$, the induced subgraph $G_{V^{\prime}}$ has vertex set $V^{\prime}$ and edge set $E_{V^{\prime}}\coloneqq E(G_{V^{\prime}})\equiv\{\{u,v\}\in E\mid\forall\,u,v\in V^{\prime}\}$. A subset of vertices $C\subseteq V$ is said to be a clique if $G_{C}$ is complete, where a complete graph is simply a graph with every pair of vertices connected by an edge.
@@
The remainder of our paper is structured as follows. Section 2 introduces and reviews several concepts central to the paper, including UU/$V$-statistics and tensors (Section 2.1), the Einstein summation (Section 2.2), and notions (including vertex elimination, treewidth, and quotient graphs) from graph theory (Section 2.3). Readers who are familiar with related materials are recommended to skip the relevant subsections. In Section 3, the new algorithm for exactly computing $U$-statistics is presented. We also illustrate how the algorithm proceeds using concrete examples to facilitate understanding. We then demonstrate the utility of our algorithm using three examples in Section 4, including computing HOIF estimators, motif counts in networks (Section 4.2), and computing distance covariances (Section 4.3). In our own opinion, Section 4 is the most important part of this paper. Section 5 concludes the paper with a discussion of future directions.
@@ 2 Preliminaries
In this section, we will review concepts important for later development, including UU/$V$-statistics and tensors (Section 2.1), the Einstein summation operation (Section 2.2), and useful notions from graph theory (Section 2.3). The way that we define $U$- or $V$-statistics may look strange to statisticians, but we find such definitions more convenient to work with when combined with the Einstein summation formulation.
2.1 UU/$V$-Statistics and Tensors

### v2

@@ Abstract
Higher-order $U$-statistics abound in fields such as statistics, machine learning, and computer science, but are known to be highly time-consuming to compute in practice. Despite their widespread appearance, a comprehensive study of their computational complexity is surprisingly lacking. This paper aims to fill this gap by presenting several results related to the computational aspect of $U$-statistics. First, we derive a useful decomposition from a $m$-th order $U$-statistic to a linear combination of $V$-statistics with orders not exceeding $m$, which are generally more feasible to compute. Second, we explore the connection between exactly computing $V$-statistics and Einstein summation, a tool often used in computational mathematics and quantum computing to accelerate tensor computations. Third, we provide an optimistic estimate of the time complexity for exactly computing $U$-statistics, based on the treewidth of a particular graph associated with the $U$-statistic kernel. The above ingredients lead to (1) a new, much more runtime-efficient algorithm to exactly compute general higher-order $U$-statistics, and (2) a more streamlined characterization of runtime complexity of computing $U$-statistics. We develop an accompanying open-source package called u-stats in both Python and R. We demonstrate through three examples in statistics that u-stats achieves impressive runtime performance compared to existing benchmarks. This paper also aspires to achieve two goals: (1) to capture the interest of researchers in both statistics and other related areas to further advance the algorithmic development of $U$-statistics and (2) to lift the burden of implementing higher-order $U$-statistics from practitioners.
Keywords: $U$-statistics, Statistical Computing, Tensors, Einstein Summation, Graphs
@@ 1 Introduction
Higher-order $U$-statistics are prevalent in various disciplines, including statistics (robins2008higher; bonvini2022fast; bonvini2024doubly; yao2018testing; breunig2024adaptive; shao2025u), machine learning (clemenccon2008ranking; clemenccon2013maximal), random graphs (chatterjee2024higher), and Bayesian networks (shpitser2011efficient), just to name a few. It is well-known that $U$-statistics are highly time-consuming to compute at large. In the worst-case, an $m$-th order $U$-statistic over a sample of size $n$ takes roughly $O(n^{m})$ steps to compute, because by definition (see Definition 1) one needs to sum over all $m$-out-of-$n$ permutations. If $n=10^{4}$ and $m=7$, $n^{m}\approx 10^{28}$!
@@
More specifically, in causal inference, robins2008higher developed a so-called “higher-order influence function (HOIF)” framework for constructing nearly rate-optimal estimators of average treatment effects under ignorability under various structural assumptions, using higher-order $U$-statistics with order roughly of $\sqrt{\log n}$. Despite their appealing statistical properties, the implementation remains a challenge in practice, especially computationally. Here we quote from wager2024causal to highlight this dilemma:
@@
One challenge with the HOIF estimator … is that to date it has been challenging to implement in practical applications …
This point is echoed in a recent review article on the future of causal inference (cinelli2025challenges) (also see zheng2025perturbed for a related discussion):
… where estimators are available that can improve on doubly robust style methods … the (higher-order) estimators are often computationally intensive and require careful tuning. There are many opportunities to make these methods more practical, automatic, and user-friendly.
@@
Our paper was initially motivated by (at least partially) addressing the computational aspect of this challenge, but we will present a general-purpose algorithm (and some theory) for computing generic higher-order $U$-statistics beyond those for HOIF estimators. We will revisit HOIF estimators in Section 4.1.
@@
To tackle the difficulty of computing higher-order $U$-statistics, statisticians and computer scientists have long resorted to the so-called “incomplete $U$-statistics”, where one sums over only a (random) subset of all $m$-out-of-$n$ permutations (blom1976some; nowicki1988subgraph; shao2025u). Another natural approach is to choose a subsample of size $n^{\prime}<n$ from the whole sample and then compute a complete $U$-statistic from that subsample. The randomization approach injects extra randomness into the final statistical results, which can produce inconsistent statistical conclusions between different randomly selected permutations/subsamples (guo2025rank). For a restricted class of higher-order $U$-statistics with decomposable and permutation-invariant kernels, kong2018estimating astutely recognized that exact computation can be accomplished by repeated matrix multiplications. For asymmetric kernels, their strategy essentially computes an incomplete $U$-statistic instead. A somewhat similar strategy has also been considered in zhang2025adaptive, in which an incomplete $U$-statistic with a special kernel structure is computed more efficiently via dynamic programming. All of the above approaches may suffer from efficiency loss by not using all the available information.
[…]
@@
The main contributions of our paper can be summarized as follows.
  • We design a new method for exactly computing higher-order $U$-statistics and another related object called $V$-statistics by leveraging the Einstein summation operation and related concepts from graph theory. Using graph theory, we provide an optimistic estimate of the time complexity of exactly computing $U$-statistics (see Section 3.3). Although the idea is conceptually simple, to the best of our knowledge, it has not been thoroughly explored in the statistical literature.
  • Despite the ubiquity of $U$-statistics, a general-purpose software package to compute them is lacking. To fill in this void, we developed a Python package u-stats (also available in R), which implemented the aforementioned algorithm. We demonstrate that the runtime of u-stats achieves the state-of-the-art performance in various examples, including computing HOIF estimators, motif counts of networks, and distance covariances in independence testing problems. In fact, as will be clear in Section 4, u-stats can dramatically improve the runtime compared to popular benchmarks. u-stats currently supports both CPU and GPU infrastructures. Having access to GPU can further improve the runtime because our algorithm heavily utilizes tensor operations such as the Einstein summation.
@@
In this part, we provide an incomplete survey of the related literature on $U$-statistics and the relevant techniques/tools to develop our new algorithm.
@@
$U$-statistics are frequently found in statistics and machine learning. For example, in certain machine learning tasks (clemenccon2008ranking; clemenccon2013maximal; shen2025engression), the empirical training risk can be written as a $U$-statistic instead of a sample mean. Various (conditional) dependence measures can also be represented as $U$-statistics (yao2018testing; he2021asymptotically; shao2025u; zhang2025adaptive); see Section 4.3 for an example. As mentioned before, in causal inference, HOIF has been shown to be a theoretically appealing framework for constructing nearly rate-optimal estimators for various causal effect parameters. Estimators based on HOIF are higher-order $U$-statistics; see Section 4.1 for concrete forms. Our work is also related to the literature on probabilistic graphical models (pearl1988probabilistic; shpitser2011efficient), in which a common problem is to compute certain marginal distributions from the joint distribution, factorized according to some Bayesian network. It can be reduced to computing $U$-statistics with kernels satisfying a particular structure to be defined later.
@@
Higher-order $U$-statistics also appear very often in statistical inference over networks, often modeled as random graphs (chatterjee2024higher; jin2025counting). For instance, chatterjee2024higher estimate motif counts (such as triangles, 4-cliques, etc.) by higher-order $U$-statistics and use these estimated motif counts to test certain properties of the underlying network. Motif counts also have widespread applications in bioinformatics (wernicke2006fanmod). As will be shown in Section 4.2, using the state-of-the-art software package igraph, for a moderately dense network, computing motif counts can be quite time consuming, necessitating the development of new algorithmic tools for $U$-statistics.
@@
Since computing UU/$V$-statistic is closely related to numerical integration using Monte Carlo sampling for which tensor operations are quite useful, our work is inspired by related techniques developed in the computational mathematics and physics communities. In particular, our algorithm utilizes the Einstein summation operation, an important tool for tensor computations. For example, in quantum computing/information, tensor networks are often used to describe many-body quantum systems and a very useful operation called tensor contraction is implemented by the Einstein summation in practice (markov2008simulating). As will be seen later, our proposed algorithm is similar to tensor contraction in spirit.
@@
In this section, we first collect some frequently used notation. We adopt the following convention: we use the demarcation $\{\}$ for a set, representing an unordered collection of distinct elements; whereas we use the demarcation $()$ for a tuple, meaning an ordered collection of elements which may contain duplicates. Given a positive integer $m$, let $m]\coloneqq\{1,2,\ldots,m\}$. To reduce clutter, we write $\bar{i}_{m}\equiv(i_{1},\cdots,i_{m})$ for short. $\mathbb{N}$ denotes the set of natural numbers that includes zero, while $\mathbb{N}^{+}$ denotes the set of all positive integers. Given a tuple $\bm{s}$, we let $\mathsf{set}(\bm{s})\coloneqq\{s_{k}\mid k\in[m]\}$ denote the set containing all the elements that appear in $\bm{s}$. Given a non-empty and finite set $S$ and a positive integer $m$, all $m$-permutations of elements in $S$ are denoted by $\mathsf{Perm}(S,m)$. All partitions of $S$ are denoted by $\Pi_{S}$. If $S$ happens to be an integer set $[m]$, $\Pi_{S}$ is simplified to $\Pi_{m}$. Given sets $S_{1},\cdots,S_{m}$, $\times_{i=1}^{m}S_{i}$ denotes their Cartesian product. When $S_{1}=\cdots=S_{m}=S$, we simply write $S^{m}\equiv\times_{i=1}^{m}S_{i}$. $S^{*}\coloneqq\cup_{m\in\mathbb{N}^{+}}S^{m}$ denotes all tuples possibly generated by $S$. Given a tuple $\bm{s}\in S^{m}$ and a tuple $\alpha\in[m]^{*}$, we let $\bm{s}[\alpha]\coloneqq(s_{\alpha_{k}})_{k=1}^{|\alpha|}$. To describe computational complexity, we write $\Theta(\cdot)$ to indicate equality up to a constant factor and $O(\cdot)$ to indicate an upper bound up to a constant factor.
@@
Terminologies from graph theory will be frequently encountered. Throughout this paper, all graphs are assumed to be finite and undirected without self-loops, but are allowed to have isolated vertices, unless stated otherwise. A graph $G$ is associated with a pair $(V,E)$, where $V\equiv V(G)$ denotes the set of vertices and $E\equiv E(G)$ denotes the set of edges in $G$. Two vertices $u$ and $v$ are said to be adjacent if $\{u,v\}\in E$. A complete graph is a graph with every pair of vertices connected by an edge, and a complete graph with $m$ vertices is denoted by $K_{m}$. For any vertex $v\in V$, its neighbor set is denoted as $N_{G}(v)\coloneqq\{u\in V\mid\{u,v\}\in E\}$, and the degree of $v$ is defined as $\deg(v,G)\coloneqq|N_{G}(v)|$. A graph $G^{\prime}$ is a subgraph of graph $G$, if $V(G^{\prime})\subseteq V(G)$ and $E(G^{\prime})\subseteq E(G)$. For any subset $V^{\prime}\subseteq V$, the induced subgraph $G_{V^{\prime}}$ has a vertex set $V^{\prime}$ and an edge set $E_{V^{\prime}}\coloneqq E(G_{V^{\prime}})\equiv\{\{u,v\}\in E\mid\forall\,u,v\in V^{\prime}\}$. A subset of vertices $C\subseteq V$ is said to be a clique if $G_{C}$ is complete.
@@
The remainder of our paper is structured as follows. Section 2 introduces and reviews several concepts central to the paper, including UU/$V$-statistics and tensors (Section 2.1), the Einstein summation (Section 2.2), and notions (including vertex elimination, treewidth, and quotient graphs) from graph theory (Section 2.3). Readers who are familiar with the related materials are recommended to skip the relevant subsections. In Section 3, the new algorithm is presented. We then demonstrate the practical merit of our algorithm using three numerical examples in Section 4, including the computations of HOIF estimators, motif counts in networks (Section 4.2), and distance covariances (Section 4.3). Section 5 concludes the paper with a discussion of future directions.
@@ 2 Preliminaries
In this section, we will review concepts important for later development, including UU/$V$-statistics and tensors (Section 2.1), the Einstein summation operation (Section 2.2), and useful notions from graph theory (Section 2.3). The way in which we define $U$- or $V$-statistics may look strange to statisticians, but we find such definitions more convenient to work with when combined with the Einstein summation formulation.
2.1 UU/$V$-Statistic and Tensor